# Nonconvergence of the Feynman–Dyson diagrammatic perturbation expansion of propagators

### v7

@@ VI Conclusions
The Feynman–Dyson diagrammatic perturbation expansions of the self-energy and Green’s function are reliably convergent and thus physically sound only in some small domains of $\omega$, which includes the central overlapping bracket around $\omega=0$ enclosing most principal roots and some low-energy satellite roots. Outside these domains, where many satellite roots reside, the perturbation expansions are frequently nonconvergent. Our mathematical analysis suggests that the convergence is assured only in an overlap of the exact and mean-field $\omega$ brackets, although the correspondences between the singularities of the exact and perturbative self-energies are hard to make except at low energies. A higher-odd-order Feynman–Dyson MBGF method can have roots that are complex or whose residues are outside the valid range of zero to one, whereas a higher-even-order Feynman–Dyson MBGF method tends to have satellite roots whose energies approach sheer HF orbital energy differences with no correlation.
@@
Infinite partial summations of diagrams can exacerbate the nonconvergence. The summation of the ladder and ring diagrams up to an infinite order by vertex renormalization (TDA) accounts for no correlation for most roots outside the central bracket. Worse still, the result of calculation changes dramatically and alternately with such a trivial computational parameter as the number of cycles taken to solve the amplitude equations in an iterative algorithm, making the method ill-defined when all roots are sought. The summation of all tower diagrams by edge renormalization (self-consistent Green’s-function method) has a factorially increasing number of roots, whose energies encroach on the frequency domains where no exact roots can be found.
@@
This is in sharp contrast with coupled-cluster theory. TDA is to IPs and EAs as LCCD is to ground-state energies since they both sum over essentially the same infinite set of ladder and ring diagrams. Despite this similarity, TDA is powerless for low- and high-lying satellite roots, while coupled-cluster theory offers the most accurate, robust, and converging approximations for all roots via EOM-CC Stanton and Bartlett (1993); Stanton and Gauss (1994); Nooijen and Bartlett (1995); Bartlett et al. (1997); Stanton and Gauss (1999); Kamiya and Hirata (2006, 2007); Hirata et al. (2000). It has been argued Nooijen and Snijders (1992, 1993); Meissner and Bartlett (1993); Kowalski et al. (2014); Peng and Kowalski (2016); Bhaskaran-Nair et al. (2016) that EOM-CC is a coupled-cluster Green’s function. The EOM formulation of MBGF Rowe (1968); Simons and Smith (1973); Herman et al. (1978, 1980a, 1980b), which also involves an infinite partial summation of diagrams, can be recast into a matrix diagonalization Baker and Pickup (1980). It may thus be free from the convergence difficulties faced by Feynman–Dyson MBGF.
@@
However, the impact of the pathologies is not limited to the Feynman–Dyson ansatz. It is hard to use higher-than-second-order self-energy and Green’s function in the ansätze that are predicated on the knowledge of all poles and residues. Such ansätze include the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995), self-consistent Green’s-function methods Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Coveney and Tew (2023), and some models of the ADC Schirmer (1982); Schirmer et al. (1983) that evaluate the static part of its self-energy by summing over all poles of a Feynman–Dyson perturbative self-energy (see Deleuze et al. Deleuze et al. (1995); Deleuze (2003) for the ADC’s lack of size- or charge-consistency, which may or may not be related to the pathologies discussed here). Insofar as the self-consistent Green’s-function methods are derived from the diagrammatic Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Gunnarsson et al. (2017); Lin and Lindsey (2018), on which the formalisms of DMFT Georges et al. (1996) are also based, the latter two may be adversely affected by the ill-defined nature of the Feynman propagator at many frequencies.
@@
The nonconvergence seems deeply rooted in the functional forms of the exact Green’s function and self-energy, featuring numerous singularities of the forms $(\omega-E_{0}+E_{I})^{-1}$ and $(\omega-E_{A}+E_{0})^{-1}$ even before a perturbation approximation is introduced. These functions are nonanalytic at many frequencies and thus cannot be expanded in converging power series, which is ultimately why perturbation theory tends to fail. This is reminiscent of the nonperturbative Bardeen–Cooper–Schrieffer theory Bardeen et al. (1957) of superconductivity, whose superconducting gap formula, $2\delta\,e^{-1/\rho V}$, is nonanalytic and cannot be expanded in a converging power series of the electron-phonon coupling $V$, which is small. To such a problem, “perturbation theory would not be easy to apply” (in page 224 of Ref. March et al. (1967)). Likewise, the Kohn–Luttinger nonconvergence Kohn and Luttinger (1960); Luttinger and Ward (1960) of finite-temperature many-body perturbation theory Jha and Hirata (2019); Hirata and Jha (2019, 2020); Jha and Hirata (2020); Hirata (2021a) is ascribed to the nonanalytic nature of the mathematical definitions of the grand potential and internal energy at $T=0$, which feature the Boltzmann factor $e^{-E/k_{\text{B}}T}$ Hirata (2021b, 2022).
In summary, the Feynman–Dyson perturbation theory is fundamentally flawed when the whole range of frequency needs to be considered, and its applicability is limited to low-energy processes. This raises a question about the robustness of the Feynman–Dyson perturbation theory as the mathematical foundation of diagrammatic or perturbative quantum field theory, which has dominated much of modern physics Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); Dyson (1993).
@@ Acknowledgements.
We thank Dr. Takashi Nakatsukasa (University of Tsukuba) for an insightful discussion on computational nuclear physics. We also thank Dr. Xiuyi Qin (CGG) for a careful reading of the manuscript and many constructive suggestions. S.H. was supported by the U.S. Department of Energy (DoE), Office of Science, Office of Basic Energy Sciences under Grant No. DE-SC0006028 and also by the Center for Scalable Predictive methods for Excitations and Correlated phenomena (SPEC), which is funded by the U.S. DoE, Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences as part of the Computational Chemical Sciences (CCS) program at Pacific Northwest National Laboratory (PNNL) under FWP 70942. PNNL is a multi-program national laboratory operated by Battelle Memorial Institute for the U.S. DoE. S.H. is a Guggenheim Fellow of the John Simon Guggenheim Memorial Foundation. I.G. was supported by Polish National Science Center under Grant No. 2020/39/O/ST4/00005 R.J.B. was supported by the Air Force Office of Scientific Research under AFOSR Award No. FA9550-19-1-0091.

### v8

@@ VI Conclusions
The Feynman–Dyson diagrammatic perturbation expansions of the self-energy and Green’s function are reliably convergent and thus physically sound only in some small domains of $\omega$, which includes the central overlapping bracket around $\omega=0$ enclosing most principal roots and some low-energy satellite roots. Outside these domains, the perturbation expansions are nonconvergent. Our mathematical analysis suggests that the convergence is assured only in an overlap of the exact and mean-field $\omega$ brackets. An odd-order MBGF($n$) method can have roots that are complex or whose residues are outside the valid range of zero to one, whereas a higher-even-order self-energy tends to consist of many vertical lines, predicting phantom poles with zero residues. A majority of satellite roots of MBGF($n$) are distributed in nearly the same frequency domains spanned by mere HF orbital energy differences and include little to no electron-correlation effects, no matter how high the perturbation order is raised.
@@
Infinite partial resummations of diagrams can exacerbate the nonconvergence. The summation of the ring and ladder diagrams up to an infinite order by vertex renormalization (TDA) deteriorates the self-energy outside the central bracket. Worse still, the result of calculation changes dramatically and alternately with the number of cycles taken to solve the amplitude equations in an iterative algorithm, making the method ill-defined when all roots are sought. The summation of all tower diagrams by edge renormalization (self-consistent Green’s function method) has a factorially increasing number of roots, whose energies encroach on the frequency domains where no exact roots can be found.
@@
This is in sharp contrast with coupled-cluster theory. TDA is to IPs and EAs as LCCD is to ground-state energies since they both sum over the same infinite set of ring and ladder diagrams. Despite this similarity, TDA is pathological for low- and high-lying satellite roots, while coupled-cluster theory via the EOM-CC formalism offers the most accurate, robust, and converging approximations for all roots obtained directly as eigenvalues of a matrix Stanton and Bartlett (1993); Stanton and Gauss (1994); Nooijen and Bartlett (1995); Bartlett et al. (1997); Stanton and Gauss (1999); Gour et al. (2005); Kamiya and Hirata (2006); Gour and Piecuch (2006); Kamiya and Hirata (2007); Hirata et al. (2000); Bartlett (2012). The EOM formulation of MBGF Rowe (1968); Simons and Smith (1973); Herman et al. (1978, 1980a, 1980b), which can also be viewed as an infinite partial resummation of diagrams, is recast into a matrix diagonalization Baker and Pickup (1980) and may thus be free from the convergence difficulties faced by MBGF($n$).
@@
The nonconvergence of the Feynman–Dyson perturbation theory, which is the mathematical foundation of quantum field theory Feynman (1948); Dyson (1949a, b); Schwinger (1951); Sakurai (1967); Dyson (1993), makes it hard to use higher-than-second-order self-energy and Green’s function in the ansätze that are predicated on the knowledge of all poles and residues. Such ansätze include the Galitskii–Migdal identity Galitskii and Migdal (1958); Koltun (1972); Holleboom and Snijders (1990); Ortiz (1995), self-consistent Green’s function methods Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Coveney and Tew (2023), and some models of the ADC Schirmer (1982); Schirmer et al. (1983) that evaluate the static part of its self-energy by summing over all poles of a Feynman–Dyson perturbative self-energy (see Deleuze et al. Deleuze et al. (1995); Deleuze (2003) for the ADC’s lack of size- or charge-consistency, which may be related to the pathologies discussed here). Insofar as the self-consistent Green’s function methods are derived from the diagrammatic Luttinger–Ward functional Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Gunnarsson et al. (2017); Lin and Lindsey (2018), on which the formalisms of DMFT Georges et al. (1996) are also based, these methods may be adversely affected by the ill-defined nature of the perturbative Feynman propagators at many frequencies.
@@
The nonconvergence is rooted in the rational-function form of the exact Green’s function, featuring numerous singularities of the forms $(\omega-E_{0}+E_{I})^{-1}$ and $(\omega-E_{A}+E_{0})^{-1}$ even before a perturbation approximation is introduced. This function is nonanalytic at many frequencies and thus cannot be expanded in converging power series, which is ultimately why perturbation theory tends to fail. This is reminiscent of the Bardeen–Cooper–Schrieffer theory of superconductivity Bardeen et al. (1957), whose superconducting gap formula, $2\delta\,e^{-1/\rho V}$, is nonanalytic and cannot be expanded in a converging power series of the electron-phonon coupling $V$, even though it is small. To such a problem, “perturbation theory would not be easy to apply” (in page 224 of Ref. March et al. (1967)). Likewise, the Kohn–Luttinger nonconvergence Kohn and Luttinger (1960); Luttinger and Ward (1960) of finite-temperature many-body perturbation theory Jha and Hirata (2019); Hirata and Jha (2019, 2020); Jha and Hirata (2020); Hirata (2021a) is ascribed to the nonanalyticity of the rational-function forms of the exact grand potential and internal energy at $T=0$ Hirata (2021b, 2022).
A Padé approximant is a power-series expansion of a rational function and is particularly promising for approximating a Feynman propagator, which is a nonanalytic rational function. Padé approximants were shown to accelerate the convergence of many-body perturbation energies, sometimes transforming a divergent series into a rapidly convergent one at no additional cost Goscinski (1967); Brändas and Goscinski (1970); Goscinski and Brändas (1971); Bartlett and Brändas (1972, 1973); Laidig et al. (1985); Hirata and Bartlett (2000b). For MBGF($n$), Padé approximants exhibit remarkable performance, largely restoring the correct functional form of the self-energy and shift its poles in the right directions by the right amounts, which is consistent with our identified cause of the nonconvergence.
@@ Acknowledgements.
We thank Professor Piotr Piecuch (Michigan State University), Dr. Xiuyi Qin (CGG), Professor Takashi Nakatsukasa (University of Tsukuba), and Professor Vojtech Vlček (University of California, Santa Barbara) for insightful discussions and constructive suggestions. S.H. was supported by the U.S. Department of Energy (DoE), Office of Science, Office of Basic Energy Sciences under Grant No. DE-SC0006028 and also by the Center for Scalable Predictive methods for Excitations and Correlated phenomena (SPEC), which is funded by the U.S. DoE, Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences as part of the Computational Chemical Sciences (CCS) program at Pacific Northwest National Laboratory (PNNL) under FWP 70942. PNNL is a multi-program national laboratory operated by Battelle Memorial Institute for the U.S. DoE. S.H. is a Guggenheim Fellow of the John Simon Guggenheim Memorial Foundation. I.G. was supported by Polish National Science Center under Grant No. 2020/39/O/ST4/00005 R.J.B. was supported by the Air Force Office of Scientific Research under AFOSR Award No. FA9550-19-1-0091.